Find the de Broglie wavelength λ for an electron moving at a speed of 1.00×106m/s. (Note that this speed is low enough that the classical momentum formula p=mv is still valid.) Recall that the mass of an electron is me=9.11×10−31kg, and Planck's constant is h=6.626×10−34J⋅s.
Answer #1

λ = h/p = h/mv = (6.626*10^-34)/(9.11*10^-31*1.0*10^6) = 7.27*10^-10 m